OpenStudy (anonymous):

a coin collection made up of pennies, nickels, and quarters contains a total of 29 coins. the number of quarters is 8 less than the number of pennies. the total face value of the coins is $1.77. how many of each denomination are there?

OpenStudy (anonymous):

Do you know how to create a model for this problem?

OpenStudy (anonymous):

mm no :/

OpenStudy (ranga):

Assume there are P pennies, N nickels, Q quarters. P + N + Q = 29 Q = P - 8 .01*P + .05*N + 0.25*Q = 1.77 Solve for P, N and Q.

OpenStudy (anonymous):

so 1.77 is the answer?